# Efficient excitation transfer in an LH2-inspired nanoscale stacked ring geometry

### v1

@@ I Introduction
With the recent technological advances in position and optical control of quantum emitters even at sub-wavelength distances in a variety of platforms [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] the interest in their intrinsic collective dynamics [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42] and possible applications in quantum information [43], and quantum sensing [44] have triggered a renewed wave of theoretical modeling. In particular, the emerging manifold of sub-radiant states [45, 46, 47, 48, 49, 24, 25, 50, 32, 51, 37, 52, 6, 7, 8] has a high potential to engineer and control coherent dynamics at extremely low loss [5] and even implement nonlinear elements [53, 54, 55]. Interestingly sub-wavelength ring structures are also omnipresent in biological light-harvesting (LH) complexes of purple photosynthetic bacteria [56, 57, 58] with their precise functionality despite extensive quantum-chemical modeling not fully understood yet. In a much-simplified quantum optical analog, the presence of finite dipole moments in each bacteriochlorophyll’s (BChl’s) enables cooperative quantum activity to explain outstanding optical properties of LH complexes [20, 32, 51, 59, 53, 39, 55, 42, 60, 61]. Specifically, a model nonameric ring with a central absorber shows superior light absorption properties [55] even in the presence of phonons [62] and enables excitation transport [32] robust to noise over longer ring chains [61].
Ring-shaped light-harvesting (LH) aggregates, in particular, LH1 and LH2 complexes exhibit circular arrangements of multiple pigments with certain rotational symmetries [57, 56, 58, 63, 64]. In photosynthesis, the coupled ring geometries of LH complexes essentially result in a highly efficient light capture and extremely fast excitation transfer mechanism [65, 66, 67, 68, 69, 70, 71, 72]. A network of coupled rings facilitates a near loss-less propagation of light energy towards the reaction center [73]. In the LH2 complex nature has designed two circular stacked layers of BChls (with sparse and dense aggregation), which simultaneously exhibit intra-ring and inter-layer [74, 75] excitation transport mechanisms [76]. Understanding nature’s design principles for light harvesting [77, 78, 79] and implementing them in artificial nanoscale devices [80, 81] are of crucial importance, as this opens up possibilities for engineering realistic ring-shaped devices that inherit this high-efficiency [82, 83, 55, 84]. Such efforts also offer a promising route to optimize energy transport [85, 32, 59, 86, 87, 88, 61] through robust and resilient ring architectures, which could provide unforeseen benefits for sustainable quantum technologies [89].
@@ II Model
Table 1: Geometric parameters in use for LH2-inspired nanorings [Fig. 1(b)]. Top ring is ${R_{1}}$ ($N_{1}=9$ emitters). Bottom ring ${R_{2}}$ consists of two rings : ${R_{2_{a}}},{R_{2_{b}}}$ ($N_{2_{a(b)}}=9$ emitters). Vertical separation is $Z_{1}$. All 27 dipoles are photoactive at 800 nm and in tangential orientation.
@@
Ring separation ($Z_{1}$) $(\nu_{i})$ $(\theta_{i},\phi_{i})$
@@
Here, the on-site collective energy shifts are neglected ($\Omega_{ii}=0$), as they lead to a simple renormalization of the bare transition frequency $\omega_{0}$. The $C_{N}$-symmetry of the ring geometry bles a Bloch eigenmode description with angular momentum quantum numbers $m$ [32]. For the geometry previously described and also depicted in Fig. 1(b), we can re-express Eq.(3) in the following form (see Appendix-A for details),
@@ III Optimizing excitation transfer between layers in an LH2 complex
LH2 Modified Nano
LH2 rings
@@ Appendix B Improved hybridization of eigenstates by altering the LH2 layer separation
From the eigenstates of one unit cell in Table-3 and 4 above, it is clear that the sub-radiant states have a higher contribution when $x=1$ (modified-geometry), than $x=2$ (biological spacings) from the bare-state $\mid 1\rangle$ [Eq.(14)], which carries the excitation of the top ring ${R_{1}}$.
@@ Appendix C Inter-layer excitation energy transfer in model nonameric LH2 rings
We theoretically consider the 3D stacked geometry of the LH2 complex of Rbl. acidophilus formed of two stacked circular layers of BChls, photoactive at 800 nm and 850 nm light, respectively [56]. We take the parameters from Ref. [96] (see Table-5) and consider each bacteriochlorophyll (BChl) as one dipole. Each dipole is a two-level emitter. In particular, top-ring ${R_{1}}$ is formed with 800 nm dipoles (sparse arrangement) and bottom-ring ${R_{2}}$ (formed of two ring ${R_{1}}$ and ${R_{2}}$) has all 850 nm dipoles (dense arrangement). The dipole moments are $|\bm{\mu}_{1}|=6.48$ D, $|\bm{\mu}_{2_{a}}|=6.41$ D and $|\bm{\mu}_{2_{b}}|=6.3$ D for rings, ${R_{1}},{R_{2_{a}}}$ and ${R_{2_{b}}}$ respectively [96]. Dipole orientations are noted in Table-5 and obtained from Ref. [96]. The total Hamiltonian for these interacting dipoles reads as,
@@
Ring radius ($r_{i}$) Ring rotation Dipole orientations
Layer separation ($Z_{1}$) $(\nu_{i})$ $(\theta_{i},\phi_{i})$
(in Å) (in $deg$) (in $deg$)

### v2

@@ I Introduction
With the recent technological advances in position and optical control of quantum emitters even at sub-wavelength distances in a variety of platforms [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], the interest in their intrinsic collective dynamics [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42] and possible applications in quantum information [43], and quantum sensing [44] have triggered a renewed wave of theoretical modeling. In particular, the emerging manifold of sub-radiant states [45, 46, 47, 48, 49, 24, 25, 50, 32, 51, 37, 52, 6, 7, 8] has a high potential to engineer and control coherent dynamics at extremely low loss [5] and even implement nonlinear elements [53, 54, 55]. Interestingly sub-wavelength ring structures are also omnipresent in biological light-harvesting (LH) complexes of purple photosynthetic bacteria [56, 57, 58] with their precise functionality despite extensive quantum-chemical modeling not fully understood yet. In a much-simplified quantum optical analog, the presence of finite dipole moments in each bacteriochlorophyll’s (BChl’s) enables cooperative quantum activity to explain outstanding optical properties of LH complexes [20, 32, 51, 59, 53, 39, 55, 42, 60, 61]. Specifically, a model nonameric ring with a central absorber shows superior light absorption properties [55] even in the presence of phonons [62] and enables excitation transport [32] robust to noise over longer ring chains [61].
Ring-shaped light-harvesting (LH) aggregates, in particular, LH1 and LH2 complexes exhibit circular arrangements of multiple pigments with certain rotational symmetries [57, 56, 58, 63, 64]. In photosynthesis, the coupled ring geometries of LH complexes essentially result in a highly efficient light capture and extremely fast excitation transfer mechanism [65, 66, 67, 68, 69, 70, 71, 72, 73]. A network of coupled rings facilitates a near loss-less propagation of light energy towards the reaction center [74]. In the LH2 complex nature has designed two circular stacked layers of BChls (with sparse and dense aggregation), which simultaneously exhibit intra-ring and inter-layer [75, 76, 77] excitation transport mechanisms [78]. Understanding nature’s design principles for light harvesting [79, 80, 81] and implementing them in artificial nanoscale devices [82, 83] are of crucial importance, as this opens up possibilities for engineering realistic ring-shaped devices that inherit this high-efficiency [84, 85, 55, 86, 87]. Such efforts also offer a promising route to optimize energy transport [88, 32, 59, 89, 90, 91, 61] through robust and resilient ring architectures, which could provide unforeseen benefits for sustainable quantum technologies [92].
@@
Following we provide a brief structure of this manuscript. In Section. II, we describe the model ring geometry under consideration and provide some theoretical descriptions together with Appendix-A. In Section. III, we theoretically explore the possibility of optimizing the inter-layer excitation transfer with model biological LH2 geometries with $C_{9}$ symmetry, when all two-level emitters are photoactive at a single wavelength (some details are in Appendix-B). We then estimate the maximum excitation energy transfer for non-identical emitter layers, as present in the biological LH2 complexes with the free-space model, and provide some discussions (some details are provided in the Appendix-C). Thereafter, we further validate our outcomes with model stacked rings with $C_{7}$, $C_{8}$ symmetry (technical details are in Appendix-D). In Section. IV, we discuss the asymmetric energy flow in the layers of the model LH2 structure. In the following Section. V, we first discuss the theoretical strategy for designing the bio-mimicked nanoscale geometry, which exhibits efficient excitation transfer via the most sub-radiant mode in between the ring layers at zero temperature (some details are provided in Appendix-E). To push our results towards a realistic experimental scenario, we discuss the effects of on-site static disorder and dephasing in the excitation transfer (also in Appendix-F). In Section. VI we supply a summary and outlook of this research endeavor. Finally in the Section. VII we provide the data availability statements.
@@ II Model
Table 1: Geometric parameters in use for LH2-inspired nanorings [Fig. 1(b)]. Top ring is ${R_{1}}$ ($N_{1}=9$ emitters). Bottom ring ${R_{2}}$ consists of two rings : ${R_{2_{a}}},{R_{2_{b}}}$ ($N_{2_{a(b)}}=9$ emitters). Vertical layer separation is $Z_{1}$. All 27 dipoles are photoactive at 800 nm and in the tangential orientation.
@@
Layer separation ($Z_{1}$) $(\nu_{i})$ $(\theta_{i},\phi_{i})$
@@
Here, the on-site collective energy shifts are neglected ($\Omega_{ii}=0$), as they lead to a simple renormalization of the bare transition frequency $\omega_{0}$. The $C_{N}$-symmetry of the ring geometry enables a Bloch eigenmode description with angular momentum quantum numbers $m$ [32]. For the geometry previously described and also depicted in Fig. 1(b), we can re-express Eq.(3) in the following form (see Appendix-A for details),
@@ III Optimizing excitation transfer between layers in an LH2 complex
Description LH2 Modified LH2 Nano rings
@@ Appendix B Improved hybridization of eigenstates by altering the LH2 layer separation
From the eigenstates of one unit cell in Table-3 and 4, it is clear that the sub-radiant states have a higher contribution when $x=1$ (modified-geometry), than $x=2$ (biological spacings) from the bare-state $\mid 1\rangle$ [Eq.(14)], which carries the excitation of the top ring ${R_{1}}$.
@@ Appendix C Inter-layer excitation energy transfer in model nonameric LH2 rings
We theoretically consider the 3D stacked geometry of the LH2 complex of Rbl. acidophilus formed of two stacked circular layers of BChls, photoactive at 800 nm and 850 nm light, respectively [56]. We take the parameters from Ref. [99] (see Table-5) and consider each bacteriochlorophyll (BChl) as one dipole. Each dipole is a two-level emitter. In particular, top-ring ${R_{1}}$ is formed with 800 nm dipoles (sparse arrangement) and bottom-ring ${R_{2}}$ (formed of two ring ${R_{1}}$ and ${R_{2}}$) has all 850 nm dipoles (dense arrangement). The dipole moments are $|\bm{\mu}_{1}|=6.48$ D, $|\bm{\mu}_{2_{a}}|=6.41$ D and $|\bm{\mu}_{2_{b}}|=6.3$ D for rings, ${R_{1}},{R_{2_{a}}}$ and ${R_{2_{b}}}$ respectively [99]. Dipole orientations are noted in Table-5 and obtained from Ref. [99].
@@
Ring size Ring rotation Dipole orientations
($r_{i}$,$Z_{1}$) (in Å) $(\nu_{i})$ (in $deg$) $(\theta_{i},\phi_{i})$ (in $deg$)
@@
The total Hamiltonian for these interacting dipoles reads as,